OpenStudy (anonymous):

Find y' sinsqrt y -3x = 2 assuming that the equation determines a differentiable function f such that y=f(x)

OpenStudy (anonymous):

you need to include brackets

OpenStudy (anonymous):

\[1/2 (\sin y)^{-1/2}(\cos y)y'-3=0\] that is what i got but i dont know how to find y' now i.e. simplify

OpenStudy (anonymous):

its not clear what the actual question is

OpenStudy (anonymous):

\[\sin \sqrt{y} - 3x=0\]

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

so sin( y^(1/2) ) -3x =2

OpenStudy (anonymous):

(1/2)y^(-1/2) dy/dx cos(y^(1/2) ) -3 =0

OpenStudy (anonymous):

or (sin y)^1/2

OpenStudy (anonymous):

\[\frac{1}{2\sqrt{y}} \frac{dy}{dx} \cos(\sqrt{y}) -3 =0 \]

OpenStudy (anonymous):

why'd we get rid of the sin? in the first term

OpenStudy (anonymous):

solve that

OpenStudy (anonymous):

why not ? lol

OpenStudy (anonymous):

derivative of sin(f (x) ) = f'(x) cos(f(x) ) yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

differentiate the inside, multiply and channge the sin to cos

OpenStudy (anonymous):

change

OpenStudy (anonymous):

ok thank you i understand that was my fall down

OpenStudy (anonymous):

you were thinking of \[\frac{1}{2} \sin^2 (\sqrt{y} ) -3x =2 \] I think

OpenStudy (anonymous):

or something similiar